A single atom of an unknown element weighs 9 x 10 -23g. What is the atomic weight of 

the element?

 


1
Expert's answer
2021-06-07T03:20:54-0400

One atomic mass unit (amu) equals:

1 amu = 1.6605 × 10−24 g

From here, the atomic mass of the element equals:

m(element) = 9 × 10-23 g / 1.6605 × 10−24 g = 5.42 × 10 = 54.2 amu


Answer: 54.2 amu